answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

The shape of the sampling distribution is always approximately normal?

welll its quiet simple to be honest 1st ASK YOUR TEACHER! that's what they are there for

If you roll a die 60 times what is the probability you will roll a one?

You add the probabilities together. For each roll, it is 1/6; therefore, you will have 60 times 1/6, which is 60/6, which is 10. Since probability is only 1 at the most, which means that it WILL happen, I would say that the probability is 1 (or 100%).

Find the distinguishable permutations of the word Texas?

Why, sixty! (Did you know that they have taxes, in Texas?!)

The calculation is 5! = 5 X 4 X 3 X 2 X 1 = 60.

There are five choices of which letter to use in the first position, four choices for the next position, and so on.

What does P with an x through the bottom of the P mean on Christian symbols?

it is a symbol for Christianity, based of the two first letters of the word Christ in Greek. Constantine saw it in the sky before a battle and was told that by that symbol he would conquer, so he painted it on his army's shields and fought extremely well!

A bag contains 9 marbles 2 green 3 blue and 4 red what is the pobability of picking a green then a non-green?

Short answer: 14/81

There is a 2/9 chance of picking a green marble the first time.

Since this leaves 8 marbles (only one of which is green; you took the other one), that leaves an 7/8 chance of picking any other colour the second time.

To find the overall probability, you simply multiply the two values (2/9 * 7/8 = 14/81). This is simple to understand. You can only get a green marble every 2 out of 9 tries. For each of these 2 times, you have a 7/8 chance of getting a different colour the second time. So, you have a 7/8 chance every 2/9 times. Simply multiply them to get the total number of possibilities.

You roll a number cube 150 times about how many times would you expect to roll 3?

Possible outcomes of one roll = 6

Successful outcomes = 1

Probability of success on each roll = 1/6

Expectation in 150 rolls = (1/6) x (150) = 25 times

A die is rolled 5 times what is the probability of rolling exactly one 4?

Possible outcomes of rolling a die 5 times = 6^5

Number of outcomes including rolling exactly one 4: 5^5 (5*5^4)

5^5/6^5 = 3125/7776 ~= 0.4019

What is the probability of not rolling a sum of 11 with two fair dice?

Since there are two permutations of two dice that sum to 11, (5-6) and (6-5), and since there are 36 permutations of rolling two dice, the probability of not rolling a sum of 11 is 34 in 36, or 17 in 18, or about 0.9444.

Who has won playing same lotto numbers?

me!!!

each 10 times they use the same numbers!!!!

here are mine:

7 - 42 - 36 - 21 - 11 - 32 - 5

use these for the next lotto!!!

you will win £1,000,000,000,000,000

How many 4 digit EVEN numbers have all 4 digits distinct?

4 digit number having 0 at last position = 9*8*7

=504

4 digit even number having 2 or 4 or 6 or 8 at last position = 8*8*7*4

=1792

(in first position, 0 cannot be fix and in last position one of digit 2,4,6,8 will be fix,

so, number of choice left for first position is 8)

So, total 4 digit even number possible = 504 + 1792

= 2296

You randomly select one card from a 52 card deck find the probability of selecting a nine or a two?

There are 4 nines and 4 twos, eight cards with the required outcome.

So the probability of a nine or two is 8/52 = 2/13 or 15.38%

What is the probability of selecting a prime number from 1 to 20?

There are eight prime numbers between 1 and 20.
2, 3, 5, 7, 11, 13, 17, 19

If you randomly choose in number then you have an 8 in 20 chance of selecting a prime.

The probability is selecting a prime number is 8/20 or 0.4

What is the Probability of drawing two spades out of 52 cards?

The probability of drawing a spade in a standard 52 card deck is 13 in 52, or 1 in 4. The probability of drawing a second spade, assuming the first spade was not replaced back into the deck, is 12 in 51. The probability, then, of drawing two spades is the product of those two probabilities, or 12 in 204, or 1 in 17.

In the game of hearts the meaning of the queen of spades?

In the game of hearts, the meaning of the queen of spades is the equivalent of 13 hearts, or half the potential score. She is also known as "Dirty Gertie" because she can be played on someone, often at the last moment, with a low score, pushing their score (which is actually a negative score) up a significant amount, but without giving them enough to invert the score for that hand. She can also be left unplayed by the leader, who finds him/her self unable to get out of the lead, said player then getting the 13 points.

Why is it necessary to use a continuity correction when using a normal distribution to approximate a binomial distribution?

It is necessary to use a continuity correction when using a normal distribution to approximate a binomial distribution because the normal distribution contains real observations, while the binomial distribution contains integer observations.